OpenStudy (anonymous):

find the area under the cure y=sqrt(9-x) on the interval [0,5]

OpenStudy (anonymous):

\[\int\limits_{0}^{5} \sqrt(9-x)\]

OpenStudy (anonymous):

the square root should go over the whole (9-x)

OpenStudy (anonymous):

how do we find the anit derivtive of that?

OpenStudy (anonymous):

You can do this by making the algebraic substitution of u = 9 - x Then, du = -dx and you can get the anti-derivative of -u^(1/2). Be careful to adjust your integration limits for the definite integral or else once you get the integral on u, convert it back to terms of "x".

OpenStudy (anonymous):

can you show me with my number

OpenStudy (anonymous):

np, but I type slowly! Hope you're patient!

OpenStudy (anonymous):

yea i need this explained so ill wait

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

I'm going to do the anti-derivative and then you can put in the limits.\[\int\limits_{}^{}\sqrt{9 - x} dx = - \int\limits_{}^{}\sqrt{u} du\]because u = 9 - x and du = -dx ,upon taking the differentials of both sides, so, the integrand becomes:\[- (\frac{ 2 }{ 3 }) u ^{\frac{ 3 }{ 2 }}\]Now, you can substitute the expression in "x" , which is 9 - x, back into this for "u"

OpenStudy (anonymous):

To check this, you might want to take the derivative and you'll see that you end up with the original integrand.

OpenStudy (anonymous):

im lost on the du part

OpenStudy (anonymous):

You might be using different symbols and techniques than what I learned. Different authors use different symbology. The easy part is thinking of u as a function of x. Then just take the derivative of u with respect to x, which can be thought of as du/dx or u'. If you keep with du/dx, then differentials will allow that you "multiply" each side by "dx" which is just the differential of x. You probably don't use du and dx. You can also just ignore it.

OpenStudy (anonymous):

but in my case what does du equal?

OpenStudy (anonymous):

The kernel of the problem is just making the substitution of u = 9 - x You can forget the theory of differentials and just concentrate on the mechanics. And taking the derivative of what I wrote will convince you of the answer. As for what du means, it is the differential change in u. It is similar to delta u where in the definition of the derivative we have delta u over delta x as delta x goes to 0.

OpenStudy (anonymous):

In short, if you learned anti-derivatives without the differential, don't worry about it in the least. Just ignore where I wrote that out. Your author apparently doesn't use that symbol.

OpenStudy (anonymous):

okayy thanks!

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

So, with making the substitution, you get: \[- \frac{ 2 }{ 3 }(9 - x)^{\frac{ 3 }{ 2 }}\]evaluated over the same interval.

OpenStudy (anonymous):

yea i understand it from there thanks!

OpenStudy (anonymous):

np!